OpenStudy (anonymous):

Which polynomial identity will prove that 64 + 27 = 91?

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (here_to_help15):

can i be rewarded a medal?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

associative property of addition

OpenStudy (here_to_help15):

yay i will give a site ok?

OpenStudy (anonymous):

There are four mathematical properties which involve addition. The properties are the commutative, associative, additive identity and distributive properties. Commutative property: When two numbers are added, the sum is the same regardless of the order of the addends. For example 4 + 2 = 2 + 4 Associative Property: When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example (2 + 3) + 4 = 2 + (3 + 4) Additive Identity Property: The sum of any number and zero is the original number. For example 5 + 0 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3

OpenStudy (anonymous):

here are the options: Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial

OpenStudy (here_to_help15):

noooo i was going to get medal = (

OpenStudy (anonymous):

oh I think 3rd one

OpenStudy (anonymous):

hey @Here_to_Help15 before I report you, stop giving sites and expect medals

OpenStudy (anonymous):

sorry! @Here_to_Help15

OpenStudy (here_to_help15):

what the heck sure fine

OpenStudy (anonymous):

Thanks for medal btw sweety!! @rayanashef98

OpenStudy (anonymous):

thanks! @Melodysim

OpenStudy (anonymous):

@Melodysim i have another one

OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = x6 – x5– x4 + 4x3 – 12x2 + 12 ?

OpenStudy (anonymous):

please!!!!!!!!!!!!